.21 1 Use a contour to evaluate S2 d8 (?+cos 8)2
Answer #1

21 I- do 2 (2+roso) use substitution, z = i8م - e & dz=ie do do -dz Fipio we know ow that jo io COSO ete 전을 = 2 2 1 -1 -izlazwe know that //ze): 2 f(z) (2-zo)? dz C 4G) (22-3) lzj=1 dz = 2元 中-24) 2 一 Z मा 中(z) = d dzLz+2+3) 中(z)=(2+24-5)-2242+5) (z+2:: I = -hi sen $(2) (2+2-13) zdz 12l=1 = 641) (27) 643 I = 40 393 IS the required integral,

Please thumbs up ?